19) Which among the following statements is TRUE? A) Red light has a shorter wavelength than violet light. B) The wavelength of light is inversely related to its energy C) As the wavelength increases, the frequency also increases. D) As the energy increases, the frequency of radiation decreases. E) none of the above 20) What is the final volume of a balloon that was initially 500.0 mL at 25 C and was then heated to 50°C? A) 461 ml B) 193 mL C) 1.00L D) 542 mL E) none of the above
Answer #1

ans 19 only B is true

A . FALSE : violet has the highest frequency and the shortest wavelength of the visible colors of light.

B . TRUE :

E= hc/?

where:
? is the wavelength, in meters (m)
c is the speed in meters per second (m/s)

C . FALSE . Wavelength is inversely proportional to frequency

D FALSE . Energy and frequency are directly proportional.

ans 20

Using Charles Law  

Initial volume/initial temperature=final volume/final temperature
V1/T1=V2/T2


V1=500mL
V2=??
T1=25C = 273+25 = 298K
T2=50C = 273+50 = 323K

V1/T1=V2/T2
500/298 = V2/323
V2 = 541.946mL = 542mL

V2 = 542 mL ( option D)